OpenStudy (anonymous):

Which of the following sets is closed under multiplication? A. {0, 1, 2, 4} B. {0, 1} C. {1, 2} D. None of the sets are closed under multiplication.

OpenStudy (amoodarya):

0*0=0 0*1=0 1*0=0 1*1=1 so {0,1} is closed under multiplying but for example {1,2} is not 1*2=2 2*1=2 1*1=1 but 2*2=4 is not in {1,2}